Please help me to solve this: Albert is 15 years old and his dad is 48. How many years ago was his dad 4 times as old as Albert was then?
Answer by stanbon(75887)

Albert is 15 years old and his dad is 48. How many years ago was his dad 4 times as old as Albert was then?
-----------------
Equation:
48-x = 4(15-x)
----
48-x = 60-4x
3x = 12
Ans:: x = 4 years ago
------------------
Cheers,
Stan H.
----------------